# Dirac Equation and Representation Dependent Scattering Phenomena

### v2

@@ 4 Non-Unitary Representation and Quantum Interference
$\eta_{1}$ $=-i(\gamma_{5}+\gamma_{2})=\begin{pmatrix}0&0&-i&-1\\ 0&0&1&-i\\ -i&1&0&0\\ -1&-i&0&0\end{pmatrix}$, (21)
$\eta_{2}$ $=\gamma_{0}+i\gamma_{5}=\begin{pmatrix}1&0&i&0\\ 0&1&0&i\\ i&0&-1&0\\ 0&i&0&-1\end{pmatrix}$. (22)
@@
which induces a metric, $g=U^{\dagger}U\neq 1$. This metric defines a modified inner product $\langle\psi|\phi\rangle_{g}=\psi^{\dagger}g\phi$ in the $\{\xi_{1},\xi_{2}\}$ representation. Natural symmetries preserving the Clifford algebra must satisfy $U^{\dagger}gU=g$ ($g$-unitarity), which differs from standard unitarity $U^{\dagger}U=I$. Note that the Hamiltonian in this representation will be pseudo-Hermitian with respect to $g$, that is, $H^{\dagger}g=gH$, with real eigenvalues and unitary time evolution [10]. This non-unitarity has notable consequences: while the bulk dispersion relation remains the same, the boundary physics and scattering properties can differ dramatically. In particular, the probability current operator:
$j_{z}=\psi^{\dagger}(\gamma_{0}+i\gamma_{2})\psi,$ (36)
differs from the standard Dirac current $j_{D}=\psi^{\dagger}\gamma^{0}\psi$, leading to distinct boundary matching conditions and hence different transmission and reflection amplitudes.

### v3

@@ 4 Non-Unitary Representation and Quantum Interference
$\eta_{1}$ $=\frac{i}{\sqrt{2}}\ \gamma_{0}(\gamma_{5}+\gamma_{2})$ (21)
$\eta_{2}$ $=i\sqrt{2}\ (\gamma_{2}+\gamma_{0})\gamma_{5}$ (22)
@@
which induces a metric, $g=U^{\dagger}U\neq 1$. This metric defines a modified inner product $\langle\psi|\phi\rangle_{g}=\psi^{\dagger}g\phi$ in the $\{\xi_{1},\xi_{2}\}$ representation. Natural symmetries preserving the Clifford algebra must satisfy $U^{\dagger}gU=g$ ($g$-unitarity), which differs from standard unitarity $U^{\dagger}U=I$. Note that the Hamiltonian in this representation will be pseudo-Hermitian with respect to $g$, that is, $H^{\dagger}g=gH$, with real eigenvalues and unitary time evolution [10]. This non-unitarity has notable consequences: while the bulk dispersion relation remains the same, the boundary physics and scattering properties can differ dramatically. In particular, the probability current operator is hermitian, given by:
$j_{z}=i\psi^{\dagger}(\gamma_{2}\gamma_{3}+\gamma_{2})\psi.$ (36)
It differs from the standard Dirac current $j_{D}=\psi^{\dagger}\gamma^{0}\psi$, leading to distinct boundary matching conditions and hence different transmission and reflection amplitudes.